OpenStudy (anonymous):

how you work out 7c+12= -4c+78

OpenStudy (lgbasallote):

put all the c on one side first. so 7c + 4c = 78-12 11c = 66 c=6

OpenStudy (anonymous):

is that the whole promblem work out or just part?

OpenStudy (lgbasallote):

if you're looking for c then that's the whole part